Parametrize Manual

Edition : auto generated by oxgentexi on 30 March 2017

OpenXM.org

1

English manual has not yet been written.

Index 2

Index

(Index is nonexistent) (Index is nonexistent)

Short Contents

Index. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

Table of Contents

Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2